A 9 year $14,000 bond paying a coupon rate of 4.50% compounded semi-annually was purchased at 96.20. Calculate the yield at the time of purchase of the bond.
Answer #1

Face value=$14000
The bond was purchased at 96.20 means, the present value=14000*96.2%=13468
Coupon rate=4.5%
Semiannual coupon rate=4.5%/2
Semiannual coupon payment=(Semiannual coupon rate)*(Face value)=(4.5%/2)*14000=315
Time period=9 years
As the interest rate is compounded semiannually, the number of periods is 9*2=18

1 Face value 14000 2 Payment 315 3 Present value - 13468 4 Time period 18 5 Yield= 2.52% 6 Formula used: RATE(W4, W2, W3, W1)
Semiannual yield =2.52%
Annualized yield=2*2.52%=5.04%